# AccentBox: Towards High-Fidelity Zero-Shot Accent Generation

### v1

@@ Abstract
While recent Zero-Shot Text-to-Speech (ZS-TTS) models have achieved high naturalness and speaker similarity, they fall short in accent fidelity and control. To address this issue, we propose zero-shot accent generation that unifies Foreign Accent Conversion (FAC), accented TTS, and ZS-TTS, with a novel two-stage pipeline. In the first stage, we achieve state-of-the-art (SOTA) on Accent Identification (AID) with 0.56 f1 score on unseen speakers. In the second stage, we condition ZS-TTS system on the pretrained speaker-agnostic accent embeddings extracted by the AID model. The proposed system achieves higher accent fidelity on inherent/cross accent generation, and enables unseen accent generation.
@@ I-A Motivation: Accent Matters in ZS-TTS
Motivated by the poor accent generation in ZS-TTS as well as the social and moral imperative for inclusive speech technology, we take an initiative to address accent-related issues in ZS-TTS. Generating accented speech in a zero-shot manner has broad and promising applications in personalised virtual assistants [9], movie dubbing [10], CAPT [7, 8], and etc.
@@ I-B Task Definition: Zero-shot Accent Generation
Previous studies on generating accented speech can be categorised into three related tasks. 1) Foreign Accent Conversion (FAC) is a speech-to-speech task that takes source speech from a target speaker as input, and converts the L2 accent in the source speech to a target L1 accent [11]. However, FAC cannot generate accented speech for any given text or generalise to unseen accent pairs. 2) Accented TTS aims to generate accented speech with high naturalness and accent fidelity with target text, accent ID, and speaker ID as input, leveraging multi-accent front-end [12, 13], Variational Auto-Encoder (VAE) [14], Diffusion [15], phoneme- and utterance-level representation learning [16, 17, 18]. Despite these studies, accented TTS remains limited by its inability to generate speech for unseen speakers or unseen accents. 3) ZS-TTS generates speech using the voice in a speech prompt (i.e. reference speech) and target text as input. Voice information derives from either speaker embeddings extracted by a pretrained speaker verification model [3, 19] or audio/speech codecs in Large Language Modelling (LLM)-based TTS [4, 20, 21]. However, none of these studies adequately addresses accent generation, with some acknowledging poor ZS-TTS performance for accented speakers [4].
@@ I-C Research Gap: Speaker-Accent Entanglement in AID and ZS-TTS
In AID, AESRC2020 benchmark [22] has been a standard. However, this data is no longer openly available with un speaker composition. A more recent benchmark, CommonAccent [23], uses a subset of Common Voice [24], which is open-source and representative of in-the-world speech data. However, our examination of the processing scripts reveals an overlap of speakers across training/validation/testing sets. The extent to which speaker-accent entanglement impacts AID performance remains unexplored, particularly when no effort is made to separate unseen speakers for testing.
@@
In ZS-TTS, the closest to our work are Zhang et al. [25, 26]. They adapt a pretrained Tacotron 2-based [27] ZS-TTS, with accent ID as input and AID as auxiliary training objective, to perform zero-shot generation for seen accents. Apart from the limitations of accented TTS, their work: 1) uses limited TTS data to learn accent embeddings, 2) relies on pre-collected accent labels in TTS data, and 3) lacks disentanglement between accent and speaker. Another closely related work by Lyth and King [28] trains an AID to pseudo-label the data and then use pseudo-generated text descriptions of the speech to control different attributes (incl. accent) in text-guided ZS-TTS. However, their work is: 1) close-sourced, with no accent generation in its open-source reproduction, Parler-TTS [^1] https://github.com/huggingface/parler-tts, 2) unclear about how the AID is trained, susceptible to speaker-accent entanglement, 3) disregarding the continuous nature of accents with pseudo-labelled discrete accent labels as TTS input condition, and 4) unable to disentangle and separately control speaker and accent in speech generation.
@@
To adress these limitations, we first propose to obtain pretrained accent embeddings from an improved AID model with speaker-accent disentanglement, termed generalisable accent identification across speakers (GenAID). This approach offers several benefits: 1) leveraging more non-TTS data to cover more speakers and accents, 2) treating accents as continuous with varying embeddings across different utterances and speakers of the same accent label, and 3) achieving greater generalisability across speakers. We then propose to condition a pretrained YourTTS-based [19] ZS-TTS on these pretrained accent embeddings, named AccentBox. AccentBox is capable of high-fidelity zero-shot accent generation and offers several advantages: 1) leveraging continuous, speaker-agnostic GenAID embeddings, 2) capable of generating unseen accents, 3) no reliance on pre-collected accent labels in TTS data, and 4) providing separate control over speaker and accent in speech generation. Readers are highly encouraged to visit our demo page [^2] https://jzmzhong.github.io/AccentBox-ICASSP2025/ where we include audio samples for accent mismatch/hallucination in current SOTA ZS-TTS (part I) and comparison between different systems and the proposed AccentBox (part IV). To summarise, our contributions are:
@@
  • To the best of our knowledge, we are the first to 1) verify and quantify the speaker-accent entanglement issue in AID data/model, and 2) highlight the accent mismatch/hallucination issue in ZS-TTS.
@@
  • We introduce novel speaker-accent disentanglement with information bottleneck and adversarial training in AID. We propose the task zero-shot accent generation and set the first benchmark for such task, unifing FAC, accented TTS, and ZS-TTS.
@@ II-A GenAID: Generalisable Accent Identification Across Speakers
In the first stage, we aim to extract continuous and speaker-agnostic accent embeddings to represent varying accents in speech. We propose an AID model that generalises across speakers, denoted GenAID (see Fig. 1). Building upon CommonAccent [23] which finetunes XLSR [29] for AID, we propose five modifications.
@@ II-A1 Validation on Unseen Speakers
To prevent the model from overfitting on seen speakers (by memorising the speaker-to-accent mapping without learning to discriminate accents), we reprocess the data and validate the model on only unseen speakers.
@@ II-A5 Adversarial Training
$\mathcal{L}=\mathcal{L}_{\text{acc\_clf}}+\alpha\cdot\mathcal{L}_{\text{MSE}}[p(y_{spk}),\mathcal{U}(|y_{spk}|)]\text{,}\vspace{-0.6em}$ (1)
@@ II-B1 Training
Since the same text spoken by speakers of different accents exhibits distinct phonetic and prosodic variations, we condition both the Transformer Text Encoder and the Stochastic Duration Predictor on the accent embeddings learned by GenAID. Compared with YourTTS, we replace the one-hot language embeddings in input with GenAID accent embeddings, as depicted by the pretrained accent encoder (orange block) in Fig. 2.
@@ II-B2 Inference
Table II outlines the different types of inference scenarios explored in this study. All reference speech, no matter for target speaker or accent information, are from unseen speakers, adhering to the zero-shot requirement. Inherent Accent Generation examines the hypothesised higher accent fidelity brought by AccentBox; Cross Accent Generation examines the hypothesised accent control and disentanglement where speaker and accent conditions mismatch; Unseen Accent Generation explores the limits of zero-shot accent generation and tests AccentBox on unseen accents.
@@ III-A2 Systems
All five modifications introduced in Sec. II-A show improvement in performance and are accumulatively added (see systems #1-#6 in Tab. IV).
@@ III-B4 Subjective Evaluation
To holistically evaluate different aspects of generated speech, we ask listeners to compare different systems based on three metrics: i) accent similarity, ii) speaker similarity, and iii) naturalness. To fully compare all systems, we conduct ABC ranking tests (Baseline vs Accent_ID vs Proposed) for inherent accent generation and AB preference tests (Accent_ID vs Proposed) for cross accent generation. The Baseline does not take any accent information as input condition and cannot perform cross accent generation, therefore not evaluated in the latter task. All listeners are recruited through Prolific [^7] https://www.prolific.com from target accent regions. 10 listeners are recruited for each utterance. Due to budget constraints, we are only able to conduct listening tests on two accents. We choose American and Irish accents, with different data size (8.03 and 3.03 hours respectively) in the finetuning data.
@@ IV-A GenAID
Table IV shows the AID resulst of different systems. On unseen speakers which we focus on, a significant 0.15 f1 score and 0.13 accuracy improvement is achieved (#1 vs #6). The best system (#6) achieves a 0.56 AID accuracy on unseen speakers, significantly better than the 0.08 random baseline. We also reduced speaker entanglement, with smaller accuracy gaps between seen and unseen speakers (0.53 vs 0.06 by #1 vs #6), and lower SCSC (0.236 vs 0.079 by #1 vs #6). Note that high accuracy on seen speakers with a large gap to unseen speakers is not desirable, as this suggests the model is memorizing speaker-accent mappings rather than learning to discriminate accents. Of all the proposed modifications, we find information bottleneck to be the most effective. We further visualise embeddings of #1 and #6 on unseen speakers using t-SNE. The best system (#6) shows better-separated accent clusters and less speaker-accent entanglement compared to the baseline (#1) in Fig. 3.
@@ IV-B AccentBox
For the task of unseen accent generation which is significantly more difficult, requiring TTS models to generalise to unseen accents, we include generated audios in the demo page with comparison between Baseline and Proposed. We leave more systematic evaluation of such task for future work.
@@
IV-B1 Inherent Accent Generation
The Proposed system achieves higher accent similarity across both objective and subjective evaluations. It outperforms other systems, including the open-source VALL-E X, in objective evaluations, regardless of the model used to extract accent embeddings. In subjective evaluations, the Proposed system consistently surpasses both the Baseline and Accent_ID systems in generating American and Irish accents, demonstrating superior accent fidelity for inherent accent generation. For speaker similarity, while subjective evaluations favor the Proposed system, objective evaluations show lower speaker cosine similarity scores, likely due to bias in the speaker verification model towards common accents or listeners’ difficulty in distinguishing accent and speaker identity. The Proposed system also shows higher naturalness when generating American accents but receives lower preference for Irish accents, potentially due to limited Irish accent data and the system’s sensitivity to monotonic prosody in reference speech. Further research with larger, more diverse datasets and refined evaluation methods is needed to better understand these discrepancies and improve performance across accents.
@@ V Conclusions
In this work, we introduce zero-shot accent generation and a novel two-stage pipeline as a benchmark. In the first stage AID, we verify, quantify, and address speaker-accent entanglement, with SOTA performance of 0.56 f1 score in 13-accent classification on unseen speakers. In the second stage zero-shot accent generation, we highlight and address the problem of accent mismatch/hallucination in ZS-TTS, with better accent fidelity in inherent/cross accent generation while enabling unseen accent generation.

### v2

@@ Abstract
While recent Zero-Shot Text-to-Speech (ZS-TTS) models have achieved high naturalness and speaker similarity, they fall short in accent fidelity and control. To address this issue, we propose zero-shot accent generation that unifies Foreign Accent Conversion (FAC), accented TTS, and ZS-TTS, with a novel two-stage pipeline. In the first stage, we achieve state-of-the-art (SOTA) on Accent Identification (AID) with 0.56 f1 score on unseen speakers. In the second stage, we condition a ZS-TTS system on the pretrained speaker-agnostic accent embeddings extracted by the AID model. The proposed system achieves higher accent fidelity on inherent/cross accent generation, and enables unseen accent generation.
@@ I-A Motivation: Accent Matters in ZS-TTS
Motivated by the poor accent generation in ZS-TTS as well as the social and moral imperative for inclusive speech technology, we take an initiative to address accent-related issues in ZS-TTS. Generating accented speech in a zero-shot manner has broad and promising applications in personalised virtual assistants [9], movie dubbing [10], CAPT [7, 8], and so on.
@@ I-B Task Definition: Zero-shot Accent Generation
Previous studies on generating accented speech can be categorised into three related tasks: 1) Foreign Accent Conversion (FAC) is a speech-to-speech task that takes source speech from a target speaker as input, and converts the L2 accent in the source speech to a target L1 accent [11]. However, FAC cannot generate accented speech for any given text or generalise to unseen accent pairs. 2) Accented TTS aims to generate accented speech with high naturalness and accent fidelity with target text, accent ID, and speaker ID as input, leveraging multi-accent front-ends [12, 13], Variational Auto-Encoder (VAE) [14], Diffusion [15], phoneme- and utterance-level representation learning [16, 17, 18]. Despite these studies, accented TTS remains limited by its inability to generate speech for unseen speakers or unseen accents. 3) ZS-TTS generates speech using the voice in a speech prompt (i.e. reference speech) and target text as input. Voice information derives from either speaker embeddings extracted by a pretrained speaker verification model [3, 19] or audio/speech codecs in Large Language Modelling (LLM)-based TTS [4, 20, 21]. However, none of these studies adequately addresses accent generation, with some acknowledging poor ZS-TTS performance for accented speakers [4].
@@ I-C Research Gap: Speaker-Accent Entanglement in AID and ZS-TTS
In AID, the AESRC2020 benchmark [22] has been a standard. However, this data is no longer openly available. A more recent benchmark, CommonAccent [23], uses a subset of Common Voice [24], which is open-source and representative of in-the-world speech data. However, our examination of the processing scripts reveals an overlap of speakers across training/validation/testing sets. The extent to which speaker-accent entanglement impacts AID performance remains unexplored, particularly when no effort is made to separate unseen speakers for testing.
@@
In ZS-TTS, the closest to our work is Zhang et al. [25, 26]. They adapt a pretrained Tacotron 2-based [27] ZS-TTS, with accent ID as input and AID as auxiliary training objective, to perform zero-shot generation for seen accents. Apart from the limitations of accented TTS, their work: 1) uses limited TTS data to learn accent embeddings, 2) relies on pre-collected accent labels in TTS data, and 3) lacks disentanglement between accent and speaker. Another closely related work by Lyth and King [28] trains an AID model to pseudo-label the data and then uses pseudo-generated text descriptions of the speech to control different attributes (incl. accent) in text-guided ZS-TTS. However, their work is: 1) close-sourced, with no accent generation in its open-source reproduction, Parler-TTS [^1] https://github.com/huggingface/parler-tts; 2) unclear about how the AID is trained, and susceptible to speaker-accent entanglement; 3) disregarding the continuous nature of accents with pseudo-labelled discrete accent labels as TTS input condition; and 4) unable to disentangle and separately control speaker and accent in speech generation.
@@
To address these limitations, we first propose to obtain pretrained accent embeddings from an improved AID model with speaker-accent disentanglement, termed generalisable accent identification across speakers (GenAID). This approach offers several benefits: 1) leveraging more non-TTS data to cover more speakers and accents, 2) treating accents as continuous with varying embeddings across different utterances and speakers of the same accent label, and 3) achieving greater generalisability across speakers. We then propose to condition a pretrained YourTTS-based [19] ZS-TTS on these pretrained accent embeddings, named AccentBox. AccentBox is capable of high-fidelity zero-shot accent generation and offers several advantages: 1) leveraging continuous, speaker-agnostic GenAID embeddings, 2) capable of generating unseen accents, 3) no reliance on pre-collected accent labels in TTS data, and 4) providing separate control over speaker and accent in speech generation. Readers are encouraged to visit our demo page [^2] https://jzmzhong.github.io/AccentBox-ICASSP2025/ where we include audio samples for accent mismatch/hallucination in current SOTA ZS-TTS (part I) and comparison between different systems and the proposed AccentBox (part IV). To summarise, our contributions are:
@@
  • To the best of our knowledge, we are the first to 1) verify and quantify the speaker-accent entanglement issue in AID data/models, and 2) highlight the accent mismatch/hallucination issue in ZS-TTS.
@@
  • We introduce novel speaker-accent disentanglement with information bottleneck and adversarial training in AID. We propose the zero-shot accent generation task and set the first benchmark for such task, unifing FAC, accented TTS, and ZS-TTS.
@@ II-A GenAID: Generalisable Accent Identification Across Speakers
In the first stage, we aim to extract continuous and speaker-agnostic accent embeddings to represent varying accents in speech. We propose an AID model that generalises across speakers, denoted GenAID (see Fig. 1). Building upon CommonAccent [23] which finetunes XLSR [29] for AID, we propose five modifications:
@@ II-A1 Validation on Unseen Speakers
To prevent the model from overfitting to seen speakers (by memorising the speaker-to-accent mapping without learning to discriminate accents), we reprocess the data and validate the model only on unseen speakers.
@@ II-A5 Adversarial Training
$\mathcal{L}=\mathcal{L}_{\text{acc\_clf}}+\alpha\cdot\mathcal{L}_{\text{MSE}}[p(y_{spk}),\mathcal{U}(|y_{spk}|)]\text{,}\vspace{-0.7em}$ (1)
@@ II-B1 Training
Since the same text spoken by speakers of different accents exhibits distinct phonetic and prosodic variations, we condition both the Transformer Text Encoder and the Stochastic Duration Predictor on the accent embeddings learned by GenAID. Compared with YourTTS, we replace the one-hot language embeddings with GenAID accent embeddings as input, as depicted by the pretrained accent encoder (orange block) in Fig. 2.
@@ II-B2 Inference
Table II outlines the different types of inference scenarios explored in this study. All reference speech use unseen speakers as target speaker or accent information, adhering to the zero-shot requirement. Inherent Accent Generation examines the hypothesised higher accent fidelity brought by AccentBox; Cross Accent Generation examines the hypothesised accent control and disentanglement where speaker and accent conditions mismatch; Unseen Accent Generation explores the limits of zero-shot accent generation and tests AccentBox on unseen accents.
@@ III-A2 Systems
All five modifications introduced in Sec. II-A bring improvement in performance and are accumulatively added (see systems #1-#6 in Tab. IV).
@@ III-B4 Subjective Evaluation
To holistically evaluate different aspects of generated speech, we ask listeners to compare different systems based on three metrics: i) accent similarity, ii) speaker similarity, and iii) naturalness. To fully compare all systems, we conduct ABC ranking tests (Baseline vs Accent_ID vs Proposed) for inherent accent generation and AB preference tests (Accent_ID vs Proposed) for cross accent generation. The Baseline does not take any accent information as input condition and cannot perform cross accent generation, therefore it is not evaluated in the latter task. All listeners are recruited through Prolific [^7] https://www.prolific.com from target accent regions. Ten listeners are recruited for each utterance. Due to budget constraints, we are only able to conduct listening tests on two accents. We choose American and Irish accents, with different data size (8.03 and 3.03 hours respectively) in the finetuning data.
@@ IV-A GenAID
Table IV shows the AID resulst of different systems. On unseen speakers, which we focus on, a significant 0.15 f1 score and 0.13 accuracy improvement is achieved (#1 vs #6). The best system (#6) achieves a 0.56 AID accuracy on unseen speakers, significantly better than the 0.08 random baseline. We also reduced speaker entanglement, with smaller accuracy gaps between seen and unseen speakers (0.53 vs 0.06 by #1 vs #6), and lower SCSC (0.236 vs 0.079 by #1 vs #6). Note that high accuracy on seen speakers with a large gap to unseen speakers is not desirable, as this suggests the model is memorizing speaker-accent mappings rather than learning to discriminate accents. Of all the proposed modifications, we find information bottleneck to be the most effective. We further visualise embeddings of #1 and #6 on unseen speakers using t-SNE. The best system (#6) shows better-separated accent clusters and less speaker-accent entanglement compared to the baseline (#1) in Fig. 3.
@@ IV-B AccentBox
For the task of Unseen Accent Generation which is significantly more difficult, requiring TTS models to generalise to unseen accents, we include generated audios on the demo page with comparison between Baseline and Proposed. We leave more systematic evaluation of such a task for future work.
IV-B1 Inherent Accent Generation
The Proposed system achieves higher accent similarity across both objective and subjective evaluations. It outperforms other systems, including the open-source VALL-E X, in objective evaluations, regardless of the model used to extract accent embeddings. In subjective evaluations, the Proposed system consistently surpasses both the Baseline and Accent_ID systems in generating American and Irish accents, demonstrating superior accent fidelity for inherent accent generation. For speaker similarity, while subjective evaluations favour the Proposed system, objective evaluations show lower speaker cosine similarity scores, likely due to bias in the speaker verification model towards common accents or listeners’ difficulty in distinguishing accent and speaker identity. The Proposed system also shows higher naturalness when generating American accents but receives lower preference for Irish accents, potentially due to limited Irish accent data and the system’s sensitivity to monotonic prosody in reference speech. Further research with larger, more diverse datasets and refined evaluation methods is needed to better understand these discrepancies and improve performance across accents.
@@ V Conclusions
In this work, we introduce zero-shot accent generation and a novel two-stage pipeline as a benchmark. In the first stage, AID, we verify, quantify, and address speaker-accent entanglement, with SOTA performance of 0.56 f1 score in 13-accent classification on unseen speakers. In the second stage, zero-shot accent generation, we highlight and address the problem of accent mismatch/hallucination in ZS-TTS, with better accent fidelity in inherent/cross accent generation while enabling unseen accent generation. In the future, we will further experiment with L2 accent generation/conversion.
Acknowledgement
This work was supported in part by the UKRI AI Centre for Doctoral Training (CDT) in Responsible and Trustworthy in-the-world NLP (Grant EP/Y030656/1), School of Informatics, and School of Philosophy, Psychology & Language Sciences, the University of Edinburgh.